OpenStudy (anonymous):

Please help me solve this: Verify by substitution that each given function is a solution of the given differential equation. y'=3x^2; y=x^3+7

OpenStudy (anonymous):

\[ y'=3x^2\] \[\int\limits{3x^2}= x^3+C\] C here happens to be 7, I'm assuming.

OpenStudy (anonymous):

thanks!